# Alleviating Hallucinations of Large Language Models through Induced Hallucinations

### 25 Dec 2023 v1

@@ 1 Introduction
Previous work Chuang et al. (2023); Tian et al. (2023a) suggests that one possible reason for hallucination might be the pre-training objective of existing LLMs, i.e., the maximum-likelihood-based next-token prediction. This objective may cause LLMs to assign non-zero probabilities to non-factual information that occurred in the training data, or to overly rely on superficial patterns learned from the training corpus rather than memorizing real-world facts (Ji et al., 2023). Nonetheless, we generally do not change this training objective considering its many other good properties, such as simplicity and generalization ability Sutskever (2023). Other researchers argue that LLM hallucinations may stem from a lack of knowledge (Zheng et al., 2023; McKenna et al., 2023). An intuitive idea for mitigating this could be injecting more knowledge into LLMs through post-hoc supervised fine-tuning (SFT). However, recent work (Schulman, 2023; Yang et al., 2023c) also highlights that the SFT process might inadvertently encourage LLMs to hallucinate by compelling them to answer questions beyond their knowledge boundaries. Furthermore, instilling a substantial amount of new factual knowledge via SFT or continual pre-training can be challenging, as it necessitates using large-scale data for downstream tasks (Chung et al., 2022; Zhang et al., 2023b), rendering the procedure computationally infeasible for most researchers today.
@@
Considering the above factors, we design a lightweight decoding method to alleviate hallucinations of LLMs, named Induce-then-Contrast Decoding (ICD). Recently, the SuperAlignment team of OpenAI unveiled the weak-to-strong generalization phenomenon (Burns et al., 2023), suggesting that weak models have the potential to elicit the capabilities of strong models. Motivated by their findings, we first construct a factually weak LLM by inducing hallucinations from the original LLM. Then we try to eliminate the non-factual knowledge internalized in the weak model from the output space of the original model through contrastive decoding (Li et al., 2023c). We show that hallucinations can be readily induced from LLMs through slight fine-tuning or zero-shot prompting, and penalizing them can effectively guide LLMs to generate more factual content. An illustration of our method is provided in Figure 1.
@@
We evaluate the effectiveness of ICD using both discrimination-based and generation-based hallucination evaluation benchmarks. Experimental results indicate that ICD significantly improves the performance of existing LLMs. For instance, when applied to TruthfulQA (Lin et al., 2022), ICD substantially improves the truthfulness of Llama2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023), making their performance comparable to the state-of-the-art ChatGPT and GPT4, as depicted in Figure 2. Additionally, when generating texts on FactScore (Min et al., 2023), ICD enables the Llama2-7B-Chat to outperform its 70B counterpart in terms of factual precision. To gain further insights into ICD, we also conduct additional analyses, such as comparing different methods for inducing hallucinations, and verifying the effectiveness of ICD across a variety of sizes and families of LLMs.
@@ 2 Related Work
Recently, various methods have been proposed to mitigate LLM hallucinations, including but not limited to strategic selection of high-quality training data (Zhou et al., 2023; Li et al., 2023d; Tian et al., 2023b), reinforcement learning from external feedback (Lightman et al., 2023; Sun et al., 2023; Yang et al., 2023c), retrieval-augmented generation (Peng et al., 2023; Vu et al., 2023; Chern et al., 2023), and the use of model uncertainty (Manakul et al., 2023; Zhang et al., 2023a). As can be observed, existing work primarily attempts to optimize LLMs to generate fewer hallucinations, which is a challenging objective. Our ICD approach, however, reframes the problem. We first aim to create a factually weak model that resembles the original model while adept at fabricating information, then subtract its knowledge from the original model’s output space to improve the factuality. We demonstrate that it could be feasible to mislead LLMs to hallucinate via custom inducements, and treating such hallucinations as a penalty term could potentially guide LLMs to be more factual.
@@ 4 Experiments
Currently, prevalent benchmarks for diagnosing hallucinations in LLMs primarily fall into two categories: discrimination-based ones and generation-based ones. In this section, we verify the effectiveness of ICD on both kinds of benchmarks.
@@ 4.1 Experimental Setup
For discrimination-based evaluation, following previous studies Chuang et al. (2023); Li et al. (2023b), we adopt the widely-used TruthfulQA [^1] https://github.com/sylinrl/TruthfulQA Lin et al. (2022). We employ multiple-choice-based metrics of TruthfulQA, specifically MC1, MC2, and MC3 scores. MC1 assesses whether models assign the highest scores to the best answer. MC2 evaluates whether the normalized probability mass for all correct answers is greater than that of the incorrect answers. MC3 examines whether each correct answer receives higher scores than all incorrect answers.
For generation-based evaluation, we employ the FActScore benchmark [^2] https://github.com/shmsw25/FActScore (Min et al., 2023). FActScore assesses the factual precision of LLMs in biography generation by breaking down generated biographies into atomic facts and comparing them with given knowledge sources. Specifically, we report the response ratio (response), the number of atomic facts per response (# facts), and the factual precision score (score) for comparison.
@@
Our experiments are basically conducted with the Llama-2 family (Touvron et al., 2023). When using our method on TruthfulQA, we induce hallucinations by fine-tuning the base model with 10k hallucinated QA pairs taken from the HaluEval [^4] https://github.com/RUCAIBox/HaluEval dataset (Li et al., 2023a). On FActScore, we fine-tune the base model with 3.5k hallucinated biographies generated by ChatGPT. More implementation details are provided in Appendix A.
@@ 4.2 Main Results
Our ICD method significantly improves the truthfulness of LLMs on TruthfulQA.
@@
ICD also reduces hallucinations in open-ended text generation on FActScore.
@@
We further launch a pair-wise automatic evaluation in Figure 3. Specifically, we utilize GPT4 to assess three dimensions of generated biographies (see more details in Appendix B), including factuality, grammaticality, and topicality. We find that ICD significantly outperforms the baseline (i.e., greedy decoding) in factuality while maintaining grammaticality and topicality.
@@ 4.3 Attempts to Use Other Methods for Hallucination Induction
From Table 1, we observe a significant truthfulness gap between the base and chat versions of Llama2. This discrepancy may be attributed to the exhaustive SFT and RLHF processes, which take honesty as an important aspect (Ouyang et al., 2022; OpenAI, 2023). This observation motivates us to directly contrast the base and chat versions of Llama2. We find that this strategy (Before/After Alignment) can also work. Notably, the improvement surpasses that of the naive CD, which could be due to the truthfulness gap between base models and aligned models being much larger than the effect of scaling up model sizes (Cheng et al., 2023).
@@ 4.4 More Analysis
To gain more in-depth insights into ICD, we perform some further analysis, as detailed below.
@@
The impact of data size for inducing hallucinations.
135103030303040404040505050506060606070707070Data size (K)
@@
In the above experiments, all the fine-tuning data used for inducing hallucinations is automatically constructed by ChatGPT. Here, we seek to figure out whether using the real failures of LLMs could lead to better performance. To this end, we generate 1,000 open-domain questions based on Wikipedia documents and ask Llama2-7B-Chat to provide answers. Then, we employ human experts to judge whether each answer is hallucinated. This procedure yields 294 real hallucinated answers, which we then utilize for fine-tuning the model for contrast. The results are displayed in Table 5. Our findings indicate that using 294 real samples could surpass the use of 1k synthetic samples on TruthfulQA, while still lagging behind the use of 10k synthetic samples. This suggests that real data might be more effective in triggering hallucinations while increasing the volume of synthetic data could narrow this gap.
@@
Extension to more LLM backbones.
To verify the applicability of our method beyond the Llama2 family, we also apply ICD to other cutting-edge open-sourced LLMs, including Baichuan2 (Yang et al., 2023a) and Mistral Jiang et al. (2023). The experimental results presented in Table 7 indicate our method generalizes well to these backbones. Moreover, it is noteworthy that the performance improvements achieved by our method in Baichuan2 and Mistral surpass those in Llama2. As we know, these two models outperform Llama2 on the standard LLM leaderboard [^5] https://opencompass.org.cn/leaderboard-llm (Contributors, 2023). This underscores our method’s ability to more effectively harness the potential of stronger backbones.
Model TruthfulQA
MC1 MC2 MC3
Baichuan2-7B-Chat 34.93 52.14 27.19
+ ICD 45.75 65.51 39.67
(+10.82) (+13.37) (+12.48)
Mistral-7B-Instruct 39.09 55.80 28.25
+ ICD 58.53 74.73 50.38
(+19.44) (+18.93) (+22.13)
As previously discussed, our method comprises two steps: inducing and contrasting. This somewhat complex pipeline motivates us to consider: is it possible to enhance the factuality of LLMs through direct fine-tuning with a selection of factual samples? Consequently, we compare our ICD method with direct fine-tuning using 3.5k factual biographies. The results are presented in Table 8. Contrary to our anticipation, we discover that direct tuning significantly impairs the factuality of the original LLM (63.8$\rightarrow$28.7), even when the training data is indeed factual. This phenomenon is interesting, and a primary explanation could be behavior cloning (Schulman, 2023), which means that SFT instructs LLMs to answer all questions without evaluating whether these questions surpass their knowledge boundaries (Yang et al., 2023c). This is further substantiated by the sharp increase in response ratio (37.5$\rightarrow$99.5). This observation suggests that mitigating hallucination via direct fine-tuning may be more challenging than expected, necessitating more sophisticated training techniques such as DPO (Tian et al., 2023b). However, our work demonstrates that inducing hallucinations via fine-tuning is feasible, and we further leverage such failure modes as a penalty term, ultimately yielding more factual responses.
@@ 5 Conclusion
This work introduces a decoding method for mitigating hallucinations in LLMs, termed induce-then-contrast decoding (ICD). Given the challenge of directly enhancing the truthfulness of LLMs, we first induce hallucinations from LLMs, and then subtract them from the output space of the original LLMs during the decoding process. Experimental results on both discrimination-based and generation-based benchmarks demonstrate that this simple method effectively improves the factuality of LLMs. We also conduct additional analysis to delve deeper into the underlying mechanism of our method, such as comparing different hallucination induction methods, assessing the effectiveness of our methods across different model sizes and types.
@@
Limitations
We think our work has the following limitations.
Additional Computational Costs.
One potential limitation of our approach is the additional computational costs introduced by contrastive decoding, which necessitates twice the forward propagation. The latency increases by about 1.6x when employing our method. In future work, we aim to explore strategies to mitigate this side effect, such as utilizing smaller models for contrast, or only training an additional head to generate hallucinations inspired by Medusa decoding (Cai et al., 2023). Regarding the GPU memory overhead, the increase is negligible due to our use of the parameter-efficient finetuning technique, i.e., LoRA (Hu et al., 2021).
Evaluation Setting.
In this work, we only evaluate our method on two hallucination benchmarks, namely TruthfulQA and FActScore. The former focuses on question answering, while the latter focuses on biographical writing, both of which can not test the universality of our method in more open domains and general tasks. The development of convincing benchmarks and metrics for diagnosing LLM hallucinations presents a significant challenge, and we plan to evaluate our method on more recent benchmarks (Chen et al., 2023; Sadat et al., 2023; Hu et al., 2023). Furthermore, we have not launched strict regression testing, which means comprehensively assessing whether ICD affects the original capacities of LLMs.

### 11 Mar 2024 v2

@@ 1 Introduction
Previous work Chuang et al. (2023); Tian et al. (2023a) suggests that one possible reason for hallucination might be the pre-training objective of existing LLMs, i.e., the maximum-likelihood-based next-token prediction. This objective may cause LLMs to assign non-zero probabilities to non-factual information that occurred in the training data, or to overly rely on superficial patterns learned from the training corpus rather than memorizing real-world facts (Ji et al., 2023). Nonetheless, this training objective still retains many good properties, such as simplicity and generalization ability Sutskever (2023), so directly modifying it may not be worth the cost. Some other researchers argue that LLM hallucinations may stem from a lack of knowledge (Zheng et al., 2023; McKenna et al., 2023). An intuitive idea for mitigating this could be injecting more knowledge into LLMs through post-hoc supervised fine-tuning (SFT). However, recent work (Schulman, 2023; Yang et al., 2023c) also highlights that the SFT process might inadvertently encourage LLMs to hallucinate by compelling them to answer questions beyond their knowledge boundaries. Furthermore, instilling a substantial amount of new factual knowledge via SFT or continual pre-training can be challenging, as it necessitates using large-scale data for downstream tasks (Chung et al., 2022; Zhang et al., 2023b), rendering the procedure computationally infeasible for most researchers today.
@@
Considering the above difficulties of mitigating hallucinations during the pre-training and SFT stages, this work designs a decoding method to alleviate LLM hallucinations, named Induce-then-Contrast Decoding (ICD). Recently, the SuperAlignment team of OpenAI unveiled the weak-to-strong generalization phenomenon (Burns et al., 2023), suggesting that weak models have the potential to elicit the capabilities of strong models. Motivated by their findings, we first construct a factually weak LLM by inducing hallucinations from the original LLM. Then we try to eliminate the non-factual information internalized in the weak model from the output space of the original model through contrastive decoding (Li et al., 2023c). We show that hallucinations can be readily induced from LLMs through slight fine-tuning or zero-shot prompting, and penalizing them can effectively guide LLMs to generate more factual content. An illustration of our method is provided in Figure 1.
@@
We evaluate the effectiveness of ICD using both discrimination-based and generation-based hallucination evaluation benchmarks. Experimental results indicate that ICD significantly improves the performance of existing LLMs. For instance, when applied to TruthfulQA (Lin et al., 2022), ICD substantially improves the truthfulness of Llama2-7B (Touvron et al., 2023) and Mistral-7B (Jiang et al., 2023), making their performance comparable to the state-of-the-art ChatGPT and GPT4, as depicted in Figure 2. Additionally, when generating texts on FactScore (Min et al., 2023), ICD enables the Llama2-7B-Chat to outperform its 70B counterpart in terms of factual precision. Experiments on LLM benchmarks, including MMLU, ARC, and AlpacaEval2.0, demonstrate that implementing ICD does not compromise the original capacity. To gain further insights into ICD, we also conduct additional analyses, such as comparing different methods for inducing hallucinations, and verifying the effectiveness of ICD across a variety of sizes and families of LLMs. The data, code, and model are available at https://github.com/hillzhang1999/ICD.
@@ 2 Related Work
Recently, various methods have been proposed to mitigate LLM hallucinations, including but not limited to strategic selection of high-quality training data (Zhou et al., 2023; Li et al., 2023e; Tian et al., 2023b), reinforcement learning from external feedback (Lightman et al., 2023; Sun et al., 2023; Yang et al., 2023c), retrieval-augmented generation (Peng et al., 2023; Vu et al., 2023; Chern et al., 2023), and the use of model uncertainty (Manakul et al., 2023; Zhang et al., 2023a). As can be observed, existing work primarily attempts to optimize LLMs to generate fewer hallucinations, which is a challenging objective. Our ICD approach, however, reframes the problem. We first aim to create a factually weak model that resembles the original model while adept at fabricating information, then subtract its knowledge from the original model’s output space to improve the factuality. We demonstrate that it could be feasible to mislead LLMs to hallucinate via custom inducements, and treating such hallucinations as a penalty term could potentially guide LLMs to be more factual.
@@ 4 Experiments
In this section, we verify the effectiveness of ICD on both discrimination-based ones and generation-based hallucination benchmarks.
@@ 4.1 Experimental Setup
For discrimination-based evaluation, following previous studies Chuang et al. (2023); Li et al. (2023b), we adopt the widely-used TruthfulQA Lin et al. (2022). We employ multiple-choice-based metrics of TruthfulQA, specifically MC1, MC2, and MC3 scores. MC1 assesses whether models assign the highest scores to the best answer. MC2 evaluates whether the normalized probability mass for all correct answers is greater than that of the incorrect answers. MC3 examines whether each correct answer receives higher scores than all incorrect answers.
For generation-based evaluation, we employ the FActScore benchmark (Min et al., 2023). FActScore assesses the factual precision of LLMs in biography generation by breaking down generated biographies into atomic facts and comparing them with given knowledge sources. Specifically, we report the response ratio (response), the number of atomic facts per response (# facts), and the factual precision score (score) for comparison.
@@
Our experiments are basically conducted with the Llama-2 family (Touvron et al., 2023). When using our method on TruthfulQA, we induce hallucinations by fine-tuning the base model with 10k hallucinated QA pairs taken from the HaluEval dataset (Li et al., 2023a). On FActScore, we fine-tune the base model with 3.5k hallucinated biographies generated by ChatGPT. More implementation details are provided in Appendix A.
@@ 4.2 Main Results
ICD significantly improves the truthfulness of LLMs on TruthfulQA.
@@
ICD reduces hallucinations in open-ended text generation on FActScore.
@@
Model MMLU ARC AlpacaEval2.0
Llama2-7B-Chat 46.35 66.41 4.91
+ ICD 46.02 67.29 5.17
Table 3: Performance before/after applying ICD on standard benchmark for evaluating the capacity of LLMs.
ICD does not hurt the original capacity.
While ICD enhances the factuality of LLMs, it is crucial to ensure that its application does not compromise the fundamental capabilities of LLMs. To verify this, we evaluate the performance of Llama2-7B-Chat before and after applying ICD on several standard LLM benchmarks, including MMLU Hendrycks et al. (2020), ARC Clark et al. (2018), and AlpacaEval2.0 Li et al. (2023d). We report 5-shot results for MMLU and ARC, and win rate compared to GPT-4-turbo outputs evaluated by GPT-4-turbo on AlpacaEval2.0. As depicted in Table 3, the incorporation of ICD effectively maintains the capacity of the LLM, which may encourage users to trustingly use ICD.
We also launch a pair-wise automatic evaluation in Figure 3. Specifically, we utilize GPT4 to assess three dimensions of generated biographies (see more details in Appendix B), including factuality, grammaticality, and topicality. We find that ICD significantly outperforms the baseline (i.e., greedy decoding) in factuality while maintaining grammaticality and topicality.
@@ 4.3 Attempts to Use Other Methods for Hallucination Induction
From Table 1, we observe a significant truthfulness gap between the base and chat versions of Llama2. This discrepancy may be attributed to the exhaustive SFT and RLHF processes, which take honesty as an important aspect (Ouyang et al., 2022; OpenAI, 2023). This observation motivates us to directly contrast the base and chat versions of Llama2. We find this strategy (Before/After Alignment) also works. Notably, the improvement surpasses that of the naive CD, which could be due to the truthfulness gap between base and aligned models being much larger than the effect of scaling up model sizes (Cheng et al., 2023).
@@ 4.4 More Analysis
Model TruthfulQA
MC1 MC2 MC3
Baichuan2-7B-Chat 34.93 52.14 27.19
+ ICD 45.75 65.51 39.67
(+10.82) (+13.37) (+12.48)
Mistral-7B-Instruct 39.09 55.80 28.25
+ ICD 58.53 74.73 50.38
(+19.44) (+18.93) (+22.13)
@@
In the above experiments, all the fine-tuning data used for inducing hallucinations is automatically constructed by ChatGPT. Here, we seek to figure out whether using the real failures of LLMs could lead to better performance. To this end, we generate 1,000 open-domain questions based on Wikipedia documents and ask Llama2-7B-Chat to provide answers. Then, we employ human experts to judge whether each answer is hallucinated. This procedure yields 294 real hallucinated answers, which we then utilize for fine-tuning the model for contrast. The results are displayed in Table 6. Our findings indicate that using 294 real samples could surpass the use of 1k synthetic samples on TruthfulQA, while still lagging behind the use of 10k synthetic samples. This suggests that real data might be more effective in triggering hallucinations while increasing the volume of synthetic data could narrow this gap. We investigate the impact of data size in Appendix C.
@@
Extension to more LLM backbones.
To verify the applicability of our method beyond the Llama2 family, we also apply ICD to other cutting-edge open-sourced LLMs, including Baichuan2 (Yang et al., 2023a) and Mistral Jiang et al. (2023). The experimental results presented in Table 7 indicate our method generalizes well to these backbones. Moreover, it is noteworthy that the performance improvements achieved by our method in Baichuan2 and Mistral surpass those in Llama2. As we know, these two models outperform Llama2 on the standard LLM leaderboard (Contributors, 2023). This underscores our method’s ability to more effectively harness the potential of stronger backbones.
As previously discussed, our method comprises two steps: inducing and contrasting. This somewhat complex pipeline motivates us to consider: is it possible to enhance the factuality of LLMs through direct fine-tuning with a selection of factual samples? Consequently, we compare our ICD method with direct fine-tuning using 3.5k factual biographies. The results are presented in Table 9. Contrary to our anticipation, we discover that direct tuning significantly impairs the factuality of the original LLM (63.8$\rightarrow$28.7), even when the training data is indeed factual. This phenomenon is interesting, and a primary explanation could be behavior cloning (Schulman, 2023), which means that SFT instructs LLMs to answer all questions without evaluating whether these questions surpass their knowledge boundaries (Yang et al., 2023c). This is further substantiated by the sharp increase in response ratio (37.5$\rightarrow$99.5). This observation suggests that mitigating hallucination via direct fine-tuning may be more challenging than expected, necessitating more sophisticated training techniques such as DPO (Tian et al., 2023b).
@@ 5 Conclusion
We introduce a decoding method for mitigating hallucinations in LLMs, termed induce-then-contrast decoding (ICD). Given the challenge of directly enhancing the truthfulness of LLMs, we first induce hallucinations from LLMs, and then penalty them from the output space of the original LLMs during decoding. Experimental results on both discrimination-based and generation-based benchmarks show that this simple method effectively improves the factuality of LLMs.
Limitations & Future Work
We think our work has the following limitations:
  1. 1.
    Additional Computational Costs. One potential limitation of our approach is the additional computational costs introduced by contrastive decoding, which necessitates twice the forward propagation. The latency increases by about 1.6x when employing our method. In future work, we aim to explore strategies to mitigate this side effect, such as utilizing smaller models for contrast, or only training an additional head to generate hallucinations inspired by Medusa decoding (Cai et al., 2023). Regarding the GPU memory overhead, the increase is negligible due to our use of the parameter-efficient finetuning technique, i.e., LoRA (Hu et al., 2021).
  2. 2.
    Evaluation Setting. In this work, we only evaluate our method on two hallucination benchmarks, namely TruthfulQA and FActScore. The former focuses on question answering, while the latter focuses on biographical writing, both of which can not test the universality of our method in more open domains and general tasks. The development of convincing benchmarks and metrics for diagnosing LLM hallucinations presents a significant challenge, and we plan to evaluate our method on more recent benchmarks (Chen et al., 2023; Sadat et al., 2023; Hu et al., 2023; Li et al., 2024).
@@ 3. Output (b), the first output from the AI system
135103030303040404040505050506060606070707070Data size (K)
Appendix B Details about GPT4 Evaluation
Appendix C The Impact of Data Size